# Curves of genus two with maps of every degree to a fixed elliptic curve

### v1

@@ Abstract.
We show that up to isomorphism there are exactly twenty pairs $(C,E)$, where $C$ is a genus-$2$ curve over $\mathbf{C}$}, where $E$ is an elliptic curve over $\mathbf{C}$}, and where for every integer $n>1$ there is a map of degree $n$ from $C$ to $E$.
@@ 1. Introduction
Curves of genus two that have nonconstant maps to elliptic curves have been studied for nearly $200$ years, beginning with work of Legendre in 1828. Below, we will briefly review some of the work of the early researchers in the field — Legendre, Jacobi, Weierstrass, Kowalevski, Poincaré, Picard, Goursat, Briochi, and others — but for now we will simply note that the problem we consider in this paper is one that could be understood by these authors, with just a little tweaking of the terminology. Namely, we address the question of whether there exists a genus-$2$ curve $C$ over the complex numbers $\mathbf{C}$}, and an elliptic curve $E$ over $\mathbf{C}$}, such that for every $n>1$ there exists a degree-$n$ morphism from $C$ to $E$. (“Is there a hyperelliptic integral that can be reduced, via transformations of every degree $n>1$, to expressions involving the same elliptic integral?”)
@@ 20. $(3+\sqrt{-1})/3$
It turns out that each quadratic forms arising in this way that represent all the integers from $2$ to $31$ is equivalent to one of the forms $q_{1}$, $q_{2}$, $q_{3}$, and $q_{4}$ given above, and therefore the $(C,E)$ pairs that we have found satisfy the conditions of Theorem 1. It is then a simple matter to compute the data presented in Table 1, and to see that there are only $20$ such pairs.
In Section 5 we compute models for the curves $C$. We close in Section 6 by proving Proposition 2.
@@
admits a degree-$2$ map to an elliptic curve. Legendre’s examples come by taking $\lambda=-1$. Later, Königsberger [23] and Picard [30, §9] each proved that every genus-$2$ curve with a degree-$2$ map to an elliptic curve occurs in Jacobi’s family.
@@
Research in these matters has continued to this day. In more modern terminology, one can fix an integer $n>1$ and study the moduli space of triples $(C,E,\varphi)$, where $\varphi\colon C\to E$ is a map of degree $n$ from a curve of genus $2$ to an elliptic curve. (Usually one demands in addition that the map be minimal, meaning that it does not factor through an isogeny $F\to E$ of degree greater than $1$.) Some work concerns the general case (see for example [11, 21, 22]), but there is also interest in considering specific small values of $n$ and constructing more or less explicit models of the corresponding moduli space, perhaps also giving equations for the triples $(C,E,\varphi)$ themselves.
@@ 2. Consequences of the existence of a degree-$2$ map
Suppose $C$ is a genus-$2$ curve over $\mathbf{C}$} with a degree-$2$ map $\varphi$ to an elliptic curve $E$. Then the special case $N=2$ of [21, Theorem 1.5] shows that there is another elliptic curve $F$ and an isomorphism $\psi\colon E[2]\to F[2]$ such that the Jacobian of $C$ is isomorphic to the quotient of $E\times F$ by the graph of $\psi$, and such that there is a degree-$2$ map $\chi\colon C\to F$.
Furthermore, if we let $G\subset(E\times F)[2]$ be the graph of $\psi$, then there is an isogeny $\Phi\colon E\times F\to\operatorname{Jac}C$ with kernel $G$, and we have a diagram
@@
The right half of this diagram is nothing other than diagram (1), which is what we want to show exists. The uniqueness of the pair $(\chi,\psi)$ up to automorphisms of $F$ is part of [21, Theorem 1.5], and the converse follows from this as well. ∎
@@
Let $P$ be point of order $2$ on $E$. Then the image of $P$ under the map from the lower left of the diagram to the $E\times F$ on the upper right is the pair $(\alpha(P),\beta(P))$, while the image of $P$ in rightmost copy of $\operatorname{Jac}C$ is 0, because the middle map from $\operatorname{Jac}C$ to $\operatorname{Jac}C$ is multiplication by $2$. Therefore, $(\alpha(P),\beta(P))$ lies in the kernel of the isogeny $\varphi^{*}\times\chi^{*}$, which is the graph of $\psi$, and it follows that $\beta(P)=\psi(\alpha(P))$. ∎
@@ 5. Models of the curves
Given values $\Delta_{E}$, $\Delta_{F}$, $\tau$, and $\sigma$ from a row of Table 1, let $K$ be a number field in which the Hilbert class polynomials of $\Delta_{E}$ and $\Delta_{F}$ split. We can compute models over $K$ for elliptic curves $E$ and $F$ with period lattices $\Lambda_{E}$ and $\Lambda_{F}$ homothetic to $\langle 1,\tau\rangle$ and $\langle 1,\sigma\rangle$, respectively. Let $L$ be an extension of $K$ over which the $2$-torsion points of $E$ and $F$ are rational. By complex approximations we can identify the $2$-torsion points $P_{1}$, $P_{\tau}$, and $P_{1+\tau}$ of $E(L)$ corresponding to the values $1/2$, $\tau/2$, and $(1+\tau)/2$ modulo $\Lambda_{E}$, and similarly we can compute the analogously-defined $2$-torsion points $Q_{1}$, $Q_{\sigma}$, and $Q_{1+\sigma}$ on $F(L)$. Let $\psi$ be the isomorphism $E[2]\to F[2]$ that sends $P_{1}$ to $Q_{\sigma}$ and $P_{\tau}$ to $Q_{1}$. Then we can use the formulas from [15, §2] to compute a curve $C$ over $L$ that corresponds to $E$, $F$, and $\psi$ as in Proposition 3.
@@
Then one can check that for integers $a,b,c,d$, we have
@@ 6.1. The quadratic form $q_{1}$
Then $n-4d^{2}$ is a positive integer and $n-4d^{2}\equiv 2,3,5\bmod 8$; a result of Dickson [10, Theorem VI] then shows that we may write
@@ 6.2. The quadratic form $q_{2}$
Then $3n-5d^{2}$ is a positive integer and $3n-5d^{2}\equiv 1,2,5,6,7\bmod 8$. Another result of Dickson [9, Theorem 15] shows that we may write
@@ 6.3. The quadratic form $q_{3}$
Then $n-3d^{2}$ is a positive integer and $n-3d^{2}\equiv 2,3,6,7\bmod 8$. Yet another result of Dickson [10, Theorem XI] shows that we may write
@@ 6.4. The quadratic form $q_{4}$
Our proof in this case depends on the parity of $n$. First let us suppose that $n>1$ is an even integer that is not a multiple of $4$. Then $n$ is congruent to $2$ or $6$ modulo $8$, so Legendre’s three-square theorem [27, p. 378] says that we may write

### v2

@@ Abstract.
We show that up to isomorphism there are exactly twenty pairs $(C,E)$, where $C$ is a genus-$2$ curve over $\mathbf{C}$}, where $E$ is an elliptic curve over $\mathbf{C}$}, and where for every integer $n>1$ there is a map of degree $n$ from $C$ to $E$. We also show that the intersection of the Humbert surfaces $H_{n^{2}}$, where $n$ ranges from $2$ to $1811$, is empty.
@@ 1. Introduction
Curves of genus two that have nonconstant maps to elliptic curves have been studied for nearly $200$ years, beginning with work of Legendre in 1828. Below, we will briefly review some of the work of the early researchers in the field — Legendre, Jacobi, Weierstrass, Kowalevski, Poincaré, Picard, Goursat, Brioschi, and others — but for now we will simply note that the problem we consider in this paper is one that could be understood by these authors, with just a little tweaking of the terminology. Namely, we address the question of whether there exists a genus-$2$ curve $C$ over the complex numbers $\mathbf{C}$}, and an elliptic curve $E$ over $\mathbf{C}$}, such that for every $n>1$ there exists a degree-$n$ morphism from $C$ to $E$. (“Is there a hyperelliptic integral that can be reduced, via transformations of every degree $n>1$, to expressions involving the same elliptic integral?”)
@@ 20. $(3+\sqrt{-1})/3$
A map $\varphi$ from a curve $C$ to an elliptic curve $E$ is said to be minimal if it does not factor through an isogeny $F\to E$ of degree greater than $1$. We note that for our pairs $(C,E)$, for some values of $n$ there are no minimal maps $C\to E$ of degree $n$. This follows from a more general result that we prove in Section 7 by using results of Kani.
Theorem 3.
Let $C$ be a curve of genus $2$ over $\mathbf{C}$}. Then for some $n$ between $2$ and $1811$, there does not exist an elliptic curve $E$ for which there exists a minimal map of degree $n$ from $C$ to $E$.
@@
It turns out that each quadratic form arising in this way that represents all the integers from $2$ to $31$ is equivalent to one of the forms $q_{1}$, $q_{2}$, $q_{3}$, and $q_{4}$ given above, and therefore the $(C,E)$ pairs that we have found satisfy the conditions of Theorem 1. It is then a simple matter to compute the data presented in Table 1, and to see that there are only $20$ such pairs.
In Section 5 we compute models for the curves $C$. In Section 6 we prove Proposition 2, and in Section 7 we prove Theorem 3.
@@
admits a degree-$2$ map to an elliptic curve. Legendre’s examples come by taking $\lambda=-1$. Later, Königsberger [25] and Picard [32, §9] each proved that every genus-$2$ curve with a degree-$2$ map to an elliptic curve occurs in Jacobi’s family.
@@
Research in these matters has continued to this day. In more modern terminology, one can fix an integer $n>1$ and study the moduli space of triples $(C,E,\varphi)$, where $\varphi\colon C\to E$ is a map of degree $n$ from a curve of genus $2$ to an elliptic curve. (Usually one demands in addition that the map be minimal, in the sense defined above.) Some work concerns the general case (see for example [12, 22, 24]), but there is also interest in considering specific small values of $n$ and constructing more or less explicit models of the corresponding moduli space, perhaps also giving equations for the triples $(C,E,\varphi)$ themselves.
@@ 2. Consequences of the existence of a degree-$2$ map
Suppose $C$ is a genus-$2$ curve over $\mathbf{C}$} with a degree-$2$ map $\varphi$ to an elliptic curve $E$. Then the special case $N=2$ of [22, Theorem 1.5] shows that there is another elliptic curve $F$ and an isomorphism $\psi\colon E[2]\to F[2]$ such that the Jacobian of $C$ is isomorphic to the quotient of $E\times F$ by the graph of $\psi$, and such that there is a degree-$2$ map $\chi\colon C\to F$.
Furthermore, if we let $G\subset(E\times F)[2]$ be the graph of $\psi$, then the isogeny $\varphi^{*}\times\psi^{*}\colon E\times F\to\operatorname{Jac}C$ has kernel $G$, and we have a diagram
@@
The right half of this diagram is nothing other than diagram (1), which is what we want to show exists. The uniqueness of the pair $(\chi,\psi)$ up to automorphisms of $F$ is part of [22, Theorem 1.5], and the converse follows from this as well. ∎
@@
Let $P$ be a point of order $2$ on $E$. Then the image of $P$ under the map from the lower left of the diagram to the $E\times F$ on the upper right is the pair $(\alpha(P),\beta(P))$, while the image of $P$ in rightmost copy of $\operatorname{Jac}C$ is 0, because the middle map from $\operatorname{Jac}C$ to $\operatorname{Jac}C$ is multiplication by $2$. Therefore, $(\alpha(P),\beta(P))$ lies in the kernel of the isogeny $\varphi^{*}\times\chi^{*}$, which is the graph of $\psi$, and it follows that $\beta(P)=\psi(\alpha(P))$. ∎
@@ 5. Models of the curves
Given values $\Delta_{E}$, $\Delta_{F}$, $\tau$, and $\sigma$ from a row of Table 1, let $K$ be a number field in which the Hilbert class polynomials of $\Delta_{E}$ and $\Delta_{F}$ split. We can compute models over $K$ for elliptic curves $E$ and $F$ with period lattices $\Lambda_{E}$ and $\Lambda_{F}$ homothetic to $\langle 1,\tau\rangle$ and $\langle 1,\sigma\rangle$, respectively. Let $L$ be an extension of $K$ over which the $2$-torsion points of $E$ and $F$ are rational. By complex approximations we can identify the $2$-torsion points $P_{1}$, $P_{\tau}$, and $P_{1+\tau}$ of $E(L)$ corresponding to the values $1/2$, $\tau/2$, and $(1+\tau)/2$ modulo $\Lambda_{E}$, and similarly we can compute the analogously-defined $2$-torsion points $Q_{1}$, $Q_{\sigma}$, and $Q_{1+\sigma}$ on $F(L)$. Let $\psi$ be the isomorphism $E[2]\to F[2]$ that sends $P_{1}$ to $Q_{\sigma}$ and $P_{\tau}$ to $Q_{1}$. Then we can use the formulas from [16, §2] to compute a curve $C$ over $L$ that corresponds to $E$, $F$, and $\psi$ as in Proposition 4.
@@
These maps all send the point $(1,0)$ on $C$ to the identity of $E$. Then one can check that for integers $a,b,c,d$, we have
@@
Remark 17.
The code in our GitHub repository also includes similar presentations of the curves and maps for the cases $\Delta_{E}=\Delta_{F}=-20$ and $\Delta_{E}=\Delta_{F}=-36$. We hope to add more examples as time allows.
Remark 18.
The examples with $\Delta_{E}=\Delta_{F}$ (rows 13 through 20 in Table 1) have a remarkable property: Each $C$ has maps of every degree to two different (but Galois conjugate) elliptic curves. If the existence of any $(C,E)$ pairs as in Theorem 1 is surprising, then surely it is even more surprising to find curves $C$ with more than one choice for EE!
@@ 6.1. The quadratic form $q_{1}$
Then $n-4d^{2}$ is a positive integer and $n-4d^{2}\equiv 2,3,5\bmod 8$; a result of Dickson [11, Theorem VI] then shows that we may write
@@ 6.2. The quadratic form $q_{2}$
Then $3n-5d^{2}$ is a positive integer and $3n-5d^{2}\equiv 1,2,5,6,7\bmod 8$. Another result of Dickson [10, Theorem 15] shows that we may write
@@ 6.3. The quadratic form $q_{3}$
Then $n-3d^{2}$ is a positive integer and $n-3d^{2}\equiv 2,3,6,7\bmod 8$. Yet another result of Dickson [11, Theorem XI] shows that we may write
@@ 6.4. The quadratic form $q_{4}$
Our proof in this case depends on the parity of $n$. First let us suppose that $n>1$ is an even integer that is not a multiple of $4$. Then $n$ is congruent to $2$ or $6$ modulo $8$, so Legendre’s three-square theorem [29, pp. 398–399] says that we may write
@@
7. The intersection of all Humbert surfaces of square degree
In this section we prove Theorem 3. Since by definition a point in the moduli space ${\mathcal{M}}_{2}$ of genus-$2$ curves lies in the Humbert surface $H_{n^{2}}$ if and only if the curve it represents has a minimal map of degree $n$ to an elliptic curve, the theorem shows that the intersection $\cap_{n>1}H_{n^{2}}$ is empty, because the intersection of the $H_{n^{2}}$ with $1<n\leq 1811$ is empty.
Kani [23, Theorem 20] shows that to every genus-$2$ curve $C$ one can associate a positive definite quadratic form $q_{C}$ in at most $3$ variables, with integer coefficients, known as the refined Humbert invariant of $C$. The invariant $q_{C}$ has the property that $C$ has a minimal map of degree $n$ to some elliptic curve if and only if $q_{C}$ represents $n^{2}$ primitively; that is, if and only if there is an integer vector $v$ with coprime entries such that $q_{C}(v)=n^{2}$. Theorem 3 is therefore a corollary of the following proposition concerning ternary quadratic forms.
Proposition 19.
Let $q$ be a positive definite ternary quadratic form with integer coefficients that does not represent $1$. Then for some $n$ with $1<n\leq 1811$, the form $q$ does not primitively represent $n^{2}$.
Proof.
Our proof is computational. Suppose there were a positive definite integer ternary form that does not represent $1$ but that primitively represents all values $n^{2}$ with $1<n\leq 1811$. We write $q$ in Minkowski-reduced form as $ax^{2}+by^{2}+cz^{2}+2rxy+2sxz+2tyz$. (Note that $r$, $s$, and $t$ may be half-integers.) By replacing some of the variables with their negations, we can assume that $r\geq 0$ and $s\geq 0$, and if $rs=0$ then we may assume that $t\geq 0$. Since the form is reduced, from [8, Lemma 1.2, p. 257] we also have
$2r$ $\leq a$ $2t$ $\leq b$ $2t$ $\geq-b$
$2s$ $\leq a$ $2t$ $\leq a+b+2r-2s$ $2t$ $\geq 2r+2s-a-b.$
$2t$ $\leq a+b-2r+2s$
We know from [36, Satz 7, p. 281] that the sequence $(a,b,c)$ is the sequence of successive minima for $q$. Since $q$ represents $4$ but does not represent $1$, we must have $2\leq a\leq 4$. If $a=2$ or $a=3$, then $ax^{2}$ does not represent $4$, so we must have $b\leq 4$ as well. On the other hand, if $a=4$, then $ax^{2}$ does not represent $9$, so $b\leq 9$.
Given the values of $a$, $b$, and $r$, consider the smallest integer $m>1$ for which $ax^{2}+by^{2}+2rxy$ does not primitively represent $m^{2}$. Suppose this $m$ is less than $1811$. (In fact, for the values of $a$, $b$, and $r$ we are considering, we will have $m\leq 5$.) Since we are assuming that $q$ primitively represents $m^{2}$, the third succesive minimum of $q$ must be at most $m^{2}$; in other words, $c\leq m^{2}$.
These observations give us an algorithm for proving the proposition. For each of the finitely many triples $(a,b,r)$ satisfying the inequalities above, we compute the value of $m$, and then let $c$ range from $b$ to $m^{2}$. Then we let $s$ and $t$ range over the values allowed by the inequalities above. For each resulting ternary form $q$, we check to see whether it primitively represents $n^{2}$ for all $n$ with $2\leq n\leq 1811$.
Lemma 20, below, gives a quick method that often can find an $n$ that is not primitively represented by $q$. If this method does not provide such an $n$ that is at most $1811$, we simply have Magma enumerate all vectors $v$ such that $q(v)\leq 49^{2}$, and look for values of $n\leq 49$ such that no primitive $v$ has $q(v)=n^{2}$. (The value $49$ is the smallest that works for all of the $q$ we must consider.)
Magma code to execute this computation is included in the GitHub repository mentioned in Section 4. We find that there are no integer ternary quadratic forms that do not represent $1$ but that do primitively represent all of the other squares less than or equal to $1811^{2}$. ∎
We note that the proposition is sharp: The form $q\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}4x^{2}+9y^{2}+15z^{2}+2xy+3xz+8yz$ primitively represents $n^{2}$ for all $n$ with $1<n<1811$ (and we give such representations in our GitHub repository), but it does not represent $1$.
Lemma 20.
Let $q=ax^{2}+by^{2}+cz^{2}+2rxy+2sxz+2tyz$ be a quadratic form with integer coefficients, let $M$ be the matrix
$\begin{pmatrix}a&r&s\\ r&b&t\\ s&t&c\end{pmatrix}$,
and set $D\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\det M$. Suppose $p$ is an odd prime such that $D$ has $p$-adic valuation $1$, such that $p$ does not divide $a$, and such that $r^{2}-ab$ is a nonsquare modulo $p$. Then $q$ does not primitively represent $p^{2}$.
Proof.
To obtain a contradiction, suppose $q$ does primitively represent $p^{2}$, and say that $x$, $y$, and $z$ are coprime integers with $p^{2}=q(x,y,z)$. We note that
(9) $q(x,y,z)=a\Bigl(x+\frac{r}{a}y+\frac{s}{a}z\Bigr)^{2}+\Bigl(\frac{ab-r^{2}}{a}\Bigr)\Bigl(y+\frac{at-rs}{ab-r^{2}}z\Bigr)^{2}+\Bigl(\frac{\det M}{ab-r^{2}}\Bigr)z^{2},$
and the coefficient of $z^{2}$ is 0 modulo $p$ by assumption. If we set $u=x+ry/a+sz/a$ and $v=y+(at-rs)z/(ab-r^{2})$, then modulo $p$ we have $0\equiv au^{2}+(ab-r^{2})v^{2}/a$. If $v$ were nonzero modulo $p$ we would have $r^{2}-ab\equiv a^{2}u^{2}/v^{2}\bmod p$, contradicting our assumption that $r^{2}-ab$ is not a square modulo $p$. Therefore $v$ is 0 modulo $p$, and so is $u$. It follows that
$y\equiv\Bigl(\frac{at-rs}{r^{2}-ab}\Bigr)z\bmod p\text{\quad and\quad}x\equiv\Bigl(\frac{bs-rt}{r^{2}-ab}\Bigr)z\bmod p.$
Since $x$, $y$, and $z$ are coprime to one another, this means that $z$ must be nonzero modulo $p$.
From (9) we see that
$0\equiv au^{2}+\Bigl(\frac{ab-r^{2}}{a}\Bigr)v^{2}+\Bigl(\frac{\det M}{ab-r^{2}}\Bigr)z^{2}\bmod p^{2}.$
Since $u$ and $v$ are divisible by $p$, the first two summands are 0 modulo $p^{2}$. But since $z$ is not divisible by $p$ and $\det M$ is not divisible by $p^{2}$, the third summand is nonzero modulo $p^{2}$. This provides us with the desired contradiction. ∎